what is an ideal gas & derive the ideal equation?

Grade:12th pass

1 Answers

arun
8 years ago
An ideal gas is a theoretical gas composed of many randomly moving point particles that do not interact except when they collide elastically. The ideal gas concept is useful because it obeys the ideal gas law, a simplified equation of state, and is amenable to analysis under statistical mechanics. Ideal gas equation is derived using kinetic theory of gases. 
                                  PV= \frac{1}{3}mNv^2           (from kinetic theory of gases)
                          and     \frac{1}{2}mNv^2=\frac{3}{2}NkT       (from equipartition of energy)
                                \Rightarrow PV = NkT
                          and     Nk =nR    (here, N is number of molecule and n is number of mole)
                                    PV =nRT